OpenStudy (anonymous):

Find the linear function f(x) that satisfies f(-2) = -5, and has a graph that is perpendicular to the line -27x-9y=4

OpenStudy (tyteen4a03):

Turn -27x-9y=4 into the slope-intercept form, you get y = (-27/9)x - 4. This means the slope of -27x-9y=4 would be (-27/9) = -3. From f(-2) = -5, we know that when x = -2, y would be -5. This means this new line passes through (-2, -5). We also know that the product of slope of 2 perpendicular lines would be -1. Let the slope of the new line be m, we have -3m = -1. Solve it, and you get 1/3. Now we put the information together again with the point-slope form, we have \(y + 5 = (1/3)(x + 2)\). Turn them into the general form, and you get -x + 3y + 13 = 0.

OpenStudy (anonymous):

thank you